OpenStudy (anonymous):

what is the intersection of the equation y=3x-4 over y= -3x -4

OpenStudy (johnweldon1993):

you would set these 2 equations equal to each other :)

jimthompson5910 (jim_thompson5910):

Start with one equation and plug in the other y=3x-4 ... start with the first equation -3x - 4 = 3x - 4 ... plug in y = -3x - 4 (the second equation)

jimthompson5910 (jim_thompson5910):

So you now need to solve -3x - 4 = 3x - 4 for x

OpenStudy (johnweldon1993):

you see how y equals..2 different equations? well....logic will say that those 2 equations are also equal correct? that is why you can set them equal to each other so 3x-4 = -3x -4 now solve for x

OpenStudy (anonymous):

so would i add 3 to -3x?

OpenStudy (johnweldon1993):

you would combine the like terms....add 3x to both sides

OpenStudy (johnweldon1993):

so 3x - 4 = -3x - 4 +3x +3x now we have 6x - 4 = -4 now we 6x by itself so we...?

OpenStudy (anonymous):

subtract 6x from -4? making -2??

OpenStudy (anonymous):

or do i add 4 to -4?

OpenStudy (johnweldon1993):

you add 4 to both sides yes (you cannot combine an "x" number and a normal number!) so 6x - 4 = -4 +4 +4 6x = 0 what is 0?

OpenStudy (johnweldon1993):

woops lol...what is x :P

OpenStudy (anonymous):

x is just 6 right? since nothing can be divided by 0?

OpenStudy (johnweldon1993):

you want x to be completely by itself so you would actually divide both sides by 6 6x = 0 -- -- 6 6 x = 0

OpenStudy (johnweldon1993):

so that says the lines will intersect when x = 0...does it make sense? want to know something funny?

OpenStudy (anonymous):

sure! and thank you, yes (: that makes plenty sense!

OpenStudy (johnweldon1993):

haha and just in case you need an extra view of it... your 2 lines are y=3x-4 and y= -3x -4 they both have the same y-intercept.... if they both cross through y at -4...then that means that at x = 0 they cross....indeed because that is what we just solved for :)

OpenStudy (anonymous):

awesome (: thanks again.

OpenStudy (johnweldon1993):

anytime :)